OpenStudy (anonymous):

Part 1: Using complete sentences, explain which method you would use to solve the following system of equations and why you chose that method. (3 points) 5x + 3y = –7 4x – 2y = –10 Part 2: Show your work and provide the solution to the system. (4 points)

jhonyy9 (jhonyy9):

- so elimination method --- multiplie first equation by 2 and second by 3 so than will get 10x +6y=-14 12x -6y = -30 -------------- assum these equations 22x 0 = -44 x= -44/22 = -2 x=-2 so than this result substitute in place of x and than will get 5*(-2) +3y = -7 -10 +3y = -7 3y= 3 y=3/3 y=1 so than x=-2 y= 1